A 300 g bird flying along at 6.3m/s sees a 10 g insect heading straight toward it with a speed of 32m/s (as measured by an observer on the ground, not by the bird). The bird opens its mouth wide and enjoys a nice lunch.

Part A What is the bird's speed immediately after swallowing? Express your answer using two significant figures.
Answer #1

Part A

From conservation of momentum,

m_{bird}v_{bird}+m_{insect}v_{insect}=(m_{bird}+m_{insect})v_{final}

\Rightarrow 300\times 6.3+10\times (-32)=(300+10)v_{final}

\Rightarrow \mathbf{v_{final}=5.06\; m/s}
Answer #2

Momentum is conserved.

m1 v1 + m2 v2 = (m1 + m2) * v

300 * 6.3 + 10 * (-32) = (300 + 10) * v

v = 5.06 m/s